# Elliptic butterflies

### v1

@@ Abstract.
We study natural evaluation and interpolation problems for elliptic functions and prove that they allow a recursive treatment using a variant of classical butterflies first introduced by Gauss. We deduce the existence of straight-line programs with $d\log(d)$ complexity for these problems and present applications to finite field arithmetic, coding theory and cryptography.
@@
Butterflyfish (freely adapted from [Sha24])
@@ 1. Introduction
One looks for a straight-line program that solves the evaluation problem using additions and multiplications by constants in $\mathbf{K}$}. The complexity of evaluation is then the minimum number of operations in such a straight-line program. One defines the complexity of interpolation in the same way.
@@
This is achieved using the well-known fast Fourier transform or FFT, first discovered by Gauss. See [BCS97, Theorem 2.6] and the historical account in [HJB85].
Bostan and Schost give in [BS05] a range of nice upper bounds for the complexity of evaluation and interpolation when the basis is the monomial basis or the Newton basis, and the evaluation set is an arithmetic sequence, a geometric sequence, or any set. They make no assumption on $\mathbf{K}$}. For example, without making any restriction on $\mathbf{K}$} nor on the evaluation set, the complexity of evaluation and interpolation in the monomial basis is bounded by an absolute constant times $M(d)\log(d)$ where $M(d)$ is the complexity of multiplying two polynomials of degree $\leqslant d$ with coefficients in $\mathbf{K}$}. Since the latter is bounded by a constant times $d\log(d)\log(\log(d))$ operations in $\mathbf{K}$}, according to a result of Schönhage and Strassen [BCS97, Theorem 2.13], we end up with a general bound of a constant times
@@
Comparing Equations (3) and (2) we see that we save a factor $\log(d)\log(\log(d))$ when $d$ is a power of two, $\mathbf{K}$} contains a primitive $d$-th root of unity, $\omega_{d}$, and the evaluation set is the set of all $d$-th roots of $1$. This gain is not marginal as is demonstrated by the many implementations and applications of FFT. Indeed general evaluation and interpolation methods often use specific ones as subroutines.
@@
To conclude this introduction we recall that there is a long tradition of interpolating on algebraic curves especially over finite fields. Among the major results, one may mention linear upper bounds on the bilinear complexity of multiplication in finite field extensions by Chudnovsky and Chudnovsky [CC88], Shparlinski, Tsfasman, Vlăduţ [STV92], Shokrollahi [Sho92], Ballet and Rolland [BR04, Bal99], Chaumine [Cha08], Randriambololona [Ran12] and others. Another major achievement was the construction of excellent codes by Goppa, Tsfasman, Vlăduţ, Zink, Ihara, García, Stichtenoth [Gop81, Gop82, TVZ82, GS95, Iha81]. Chudnovsky and Chudnovsky express in [CC89, Section 6] the intuition that elliptic curves with smooth order could be of some use to accelerate some specific interpolation problems in genus 0. Using similar ideas, Ben-Sasson, Carmon, Kopparty, and Levit show in [Ben+23] how to quickly compute the values of a polynomial at a set $S^{\prime}$ from its value at another set $S$ when $S$ and $S^{\prime}$ are specific subsets of finite fields i.e. $x$-coordinates of well chosen points on a well chosen elliptic curve. Starting from Chudnovsky’s intuition we prefer to stay in the elliptic world where all these ideas and objects originate from, and develop an explicit analogue of butterflies in this context. We also prefer to invoke evaluation and interpolation rather than Fourier transform, because in our humble opinion, a Fourier transform is something different. See e.g. [BCS97, Section 13.5] or [CG23, Section 6.1].
@@ 2. Summary and organization of the article
In this article we consider specific evaluation and interpolation problems in genus one. We consider an elliptic curve $E$ over a field $\mathbf{K}$} of odd characteristic. We let $d$ be a power of two and we assume that there is a point $t$ of order $d$ in $E({\mathbf{K}})$. We let
@@
be the degree $d$ divisor associated to the group $\langle t\rangle$} generated by $t$. We take $b$ to be a $\mathbf{K}$}-rational point on $E$ such that
$db\not=0.$
We set
@@
A first application, presented in Section 13, is the construction of normal bases for finite field extensions of degree a power of two, allowing multiplication in time ${\mathcal{Q}}d\log(d)$. We build on the construction in [CL09] of an equivariant version of Chudnovsky’s interpolation method [CC88], and we use our complexity estimates for evaluation, interpolation and reduction in this context.
In Section 14, we construct Goppa codes in genus one obtained by restricting the evaluation map $\mathop{\rm{ev}}\nolimits$ to the subspace of functions that are invariant by the elliptic involution $P\mapsto-P$ on $E$. The resulting code is a genus zero Goppa code. The advantage of seeing it as a genus one code is that we benefit from the acceleration due to the presence of the automorphism group $\langle t\rangle$}. This produces MDS $[d,d/2,d/2+1]$-codes that can be encoded and checked in time $d\log(d)$ and decoded up to half the minimum distance at the expense of ${\mathcal{Q}}d\log^{2}(d)\log(\log(d))$ operations and comparisons in $\mathbf{K}$}. The only constraint for the existence of such codes is that the size of $\mathbf{K}$} be large enough with respect to $d$. This extends Lacan and Soro’s result [SL10] for Fermat fields.
@@ 3. Algorithms and data structures
For ease of reading, our pseudo-code descriptions of these algorithms are given in a recursive form. However, like the classical butterfly formulation of FFT algorithms à la Cooley–Tukey, they can be very easily unrolled to yield iterative functions. In fact, our algorithms are all straight-line programs: apart from precomputations of constants, they contain no loops or conditionals, and consist of a sequence of operations, each applied to previously computed elements.
@@ 4. Elliptic functions
We recall a few formulae from [CL09] regarding low-degree functions on elliptic curves. We let $\mathbf{K}$} be a field and $E$ the elliptic curve over $\mathbf{K}$} defined by the Weierstrass equation
@@ 5. Bases for elliptic function field extensions
Let $\mathbf{K}$} be a field with odd characteristic and let $E$ be an elliptic curve over $\mathbf{K}$}. Let $t$ be a point of order $d=2^{\delta}$ in $E({\mathbf{K}})$. We assume that $\delta\geqslant 1$. The group $\langle t\rangle$} generated by $t$ can be seen as an effective divisor of degree $d$ on $E$. We let ${\mathcal{L}}({\langle t\rangle})$ be the associated $\mathbf{K}$}-linear space. We construct two bases for ${\mathcal{L}}({\langle t\rangle})$ consisting of functions of degree $\leqslant 2$. We start with the following lemma from [CL09, Lemma 4].
@@
We deduce from Equations (6) and (7) that base change between $(u_{l})_{l\in{\mathbf{Z}}/d{\mathbf{Z}}}$ and $(v_{l})_{l\in{\mathbf{Z}}/d{\mathbf{Z}}}$ requires no more than ${\mathcal{Q}}d$ additions and multiplications by a constant in $\mathbf{K}$}.
@@ 6. A Vélu isogeny of degree two
the point of order two in the group $\langle t\rangle$}. Following Vélu [Vél71, Vél78] we write $T=(x(T),y(T))$ and set
@@ 7. Evaluation using elliptic butterflies
Under the hypotheses and notation of Sections 5 and 6, we let $b$ be a point in $E({\mathbf{K}})$ such that
$db\not=0.$
@@
$\hat{\varphi}(d^{\prime}b^{\prime})\not=0$
@@
Parameters The constants $\vec{{\mathfrak{a}}}\,^{\prime}$, $\vec{{\mathfrak{b}}}$ and $\vec{{\mathfrak{c}}}$ are defined by ${\mathfrak{b}}_{0}={\mathfrak{b}}_{\ast}$ , ${\mathfrak{c}}_{0}={\mathfrak{c}}_{\ast}$ and by Eq. (8) and (20) . The constant $\vec{{\mathfrak{x}}}\,^{\prime}$ denotes $(x^{\prime}(b^{\prime}+l\,t^{\prime})-x^{\prime}(U^{\prime}))_{l\in{\mathbf{Z}}/d^{\prime}{\mathbf{Z}}}$ . The constant $\vec{{\mathfrak{v}}}\,^{\prime}$ denotes $(v^{\prime}_{l}(U^{\prime}))_{l\in{\mathbf{Z}}/d^{\prime}{\mathbf{Z}}}$ . The constant $\vec{{\mathfrak{t}}}$ denotes $(\theta(b+l\,t))_{l=0,\,\ldots\,d^{\prime}-1}$ . The constant $\vec{{\mathfrak{m}}}\,^{\prime}$ denotes $\vec{{\mathfrak{b}}}\,\diamond\,\vec{{\mathfrak{v}}}\,^{\prime}$ and $\vec{{\mathfrak{n}}}\,^{\prime}$ denotes $\vec{{\mathfrak{m}}}\,^{\prime}+\sigma^{-1}(\vec{{\mathfrak{c}}}\,\diamond\,\vec{{\mathfrak{v}}}\,^{\prime})$ .
@@
There exists a constant $\mathcal{Q}$} such that the following is true. Let $\mathbf{K}$} be a field with odd characteristic. Let $E$ be an elliptic curve over $\mathbf{K}$}. Let $\delta\geqslant 1$ be an integer. Let $t$ be a point of order $d=2^{\delta}$ in $E({\mathbf{K}})$. Let $b$ be a point in $E({\mathbf{K}})$ such that $db\not=0$. There exists a straight-line program that on input a function $f$ in ${\mathcal{L}}({\langle t\rangle})$ given by its coordinates in either basis $u$ or $v$ defined in Section 5, computes the values $f(b+lt)$ for $l\in[0,d-1]$ at the expense of ${\mathcal{Q}}d\log(d)$ operations in $\mathbf{K}$}.
This results from the discussion above. For clarity, the evaluation procedure is presented as concise pseudo-code in Figure 2.
$\square$
@@ 8. Interpolation
According to Equation (13) these $d^{\prime}$ scalars are the coordinates of $f^{1}=\theta\cdot f^{-}$ in the basis of ${\mathcal{L}}({\langle t^{\prime}\rangle}+O^{\prime}-U^{\prime})$ made of the $\theta\cdot(u_{l}-u_{l+d^{\prime}})$ for $l\in[0,d^{\prime}-1]$. We will use an auxiliary function $\xi_{b}$ on $E^{\prime}$ having degree $d^{\prime}+1$, polar divisor ${\langle t^{\prime}\rangle}+O^{\prime}$, and vanishing at $b^{\prime}+lt^{\prime}$ for every $l\in[0,d^{\prime}-1]$. Such a function is unique up to a multiplicative constant, and it has no pole and no zero at $U^{\prime}$ because $db\not=0$. We therefore assume that $\xi_{b}(U^{\prime})=1$. There exists $d^{\prime}+1$ scalars ${\mathfrak{d}}_{0}$, ${\mathfrak{d}}_{1}$, …, ${\mathfrak{d}}_{d^{\prime}-1}$, $\mathfrak{d}$$_{\ast}$ such that
@@
$f^{\star}=f^{\star}(U^{\prime})+\sum_{1\leqslant l\leqslant d^{\prime}-1}f^{\star}_{l}\cdot(v^{\prime}_{l}-v^{\prime}_{l}(U^{\prime}))$
@@
This system has a very special form. One says it is cyclic bidiagonal. Solving it requires ${\mathcal{Q}}d$ operations in $\mathbf{K}$} as recalled in section 9. In the special case when $d^{\prime}=1$ we simply invoke Equation (17).
@@
There exists a constant $\mathcal{Q}$} such that the following is true. Let $\mathbf{K}$} be a field with odd characteristic. Let $E$ be an elliptic curve over $\mathbf{K}$}. Let $\delta\geqslant 1$ be an integer. Let $t$ be a point of order $d=2^{\delta}$ in $E({\mathbf{K}})$. Let $b$ be a point in $E({\mathbf{K}})$ such that $db\not=0$. There exists a straight-line program that on input $d$ scalars $(\alpha_{l})_{0\leqslant l\leqslant d-1}$, computes a function $f$ in ${\mathcal{L}}({\langle t\rangle})$, given by its coordinates in either basis $u$ or $v$ defined in Section 5, such that $f(b+lt)=\alpha_{l}$ for every $l\in[0,d-1]$, at the expense of ${\mathcal{Q}}d\log(d)$ operations in $\mathbf{K}$}.
This results from the discussion above. To aid understanding, the interpolation procedure is shown in compact pseudo-code in Figure 3.
$\square$
@@
Parameters The constants $\vec{{\mathfrak{a}}}\,^{\prime}$, $\vec{{\mathfrak{b}}}$ and $\vec{{\mathfrak{c}}}$ are defined by ${\mathfrak{b}}_{0}={\mathfrak{b}}_{\ast}$ , ${\mathfrak{c}}_{0}={\mathfrak{c}}_{\ast}$ and by Eq. (8) and (20) . The constant $\vec{{\mathfrak{d}}}$ is defined by Eq. (21). The constant $\vec{{\mathfrak{v}}}\,^{\prime}$ denotes $(v^{\prime}_{l}(U^{\prime}))_{l\in{\mathbf{Z}}/d^{\prime}{\mathbf{Z}}}$ . The constant $\vec{{\mathfrak{t}}}$ denotes $(\theta(b+l\,t))_{l=0,\,\ldots\,d^{\prime}-1}$ .
@@
9. Cyclic bidiagonal linear systems
The determinant of $M$ is
$\det(M)=\prod_{i=0}^{d-1}b_{i}-(-1)^{d}\prod_{i=0}^{d-1}c_{i}.$

### v2

@@ Abstract.
We study natural evaluation and interpolation problems for elliptic functions and prove that they allow a recursive treatment using a variant of classical butterflies first introduced by Gauss. We deduce the existence of straight-line programs with complexity scaling with $d\log(d)$ for these problems and present applications to finite field arithmetic, coding theory and cryptography.
@@
Butterflyfish (freely adapted from [31])
@@ 1. Introduction
One looks for straight-line programs that solve the evaluation and interpolation problems using additions and multiplications by precomputed constants in $\mathbf{K}$}. Such programs contain no loops or conditionals, and consist of a sequence of operations involving the inputs, the constants, and previously computed elements. The complexity is then measured by the number of operations.
@@
This is achieved using the well-known fast Fourier transform or FFT, first discovered by Gauss. See [10, Theorem 2.6] and the historical account in [23].
Bostan and Schost give in [9] a range of nice upper bounds for the complexity of evaluation and interpolation when the basis is the monomial basis or the Newton basis, and the evaluation set is an arithmetic sequence, a geometric sequence, or any set. They make no assumption on $\mathbf{K}$}. For example, without making any restriction on $\mathbf{K}$} nor on the evaluation set, the complexity of evaluation and interpolation in the monomial basis is bounded by an absolute constant times $M(d)\log(d)$ where $M(d)$ is the complexity of multiplying two polynomials of degree $\leqslant d$ with coefficients in $\mathbf{K}$}. Since the latter is bounded by a constant times $d\log(d)\log(\log(d))$ operations in $\mathbf{K}$}, according to a result of Schönhage and Strassen [10, Theorem 2.13], we end up with a general bound of a constant times
@@
Comparing Equations (3) and (2) we see that we save a factor $\log(d)\log(\log(d))$ when $d$ is a power of two, $\mathbf{K}$} contains a primitive $d$-th root of unity, $\omega_{d}$, and the evaluation set is the set of all $d$-th roots of $1$. We notice that in the specific case of evaluation, resp. interpolation, in the monomial basis over a geometric progression of size $d$, Bostan and Schost [9, Prop. 5] also give an algorithm whose complexity decreases to $M(d)+O(d)$, resp. $2M(d)+O(d)$. This still differs from Equation (2) by a factor of $\log(\log(d))$. This gain is not marginal as is demonstrated by the many implementations and applications of FFT. Indeed general evaluation and interpolation methods often use specific ones as subroutines.
@@
To conclude this introduction we recall that there is a long tradition of interpolating on algebraic curves especially over finite fields. Among the major results, one may mention linear upper bounds on the bilinear complexity of multiplication in finite field extensions by Chudnovsky and Chudnovsky [12], Shparlinski, Tsfasman, Vlăduţ [34], Shokrollahi [32], Ballet and Rolland [3, 4], Chaumine [11], Randriambololona [28] and others. Another major achievement was the construction of excellent codes by Goppa, Tsfasman, Vlăduţ, Zink, Ihara, García, Stichtenoth [19, 20, 37, 18, 24]. Chudnovsky and Chudnovsky express in [13, Section 6] the intuition that elliptic curves with smooth order could be of some use to accelerate some specific interpolation problems in genus 0. Using similar ideas, Ben-Sasson, Carmon, Kopparty, and Levit show in [5] how to quickly compute the values of a polynomial at a set $S^{\prime}$ from its value at another set $S$ when $S$ and $S^{\prime}$ are specific subsets of finite fields i.e. $x$-coordinates of well chosen points on a well chosen elliptic curve. Starting from Chudnovsky’s intuition we prefer to stay in the elliptic world where all these ideas and objects originate from, and develop an explicit analogue of butterflies in this context. We also prefer to invoke evaluation and interpolation rather than Fourier transform, because in our humble opinion, a Fourier transform is something different. See e.g. [10, Section 13.5] or [14, Section 6.1].
Finally, we would like to stress that we do not address the problem of polynomial multiplication in this work. We stick to elliptic functions. In principle, one could disguise two polynomials as elliptic functions, then multiply them out using elliptic butterflies. The cost of translation, however, would be $O(d\log^{2}(d))$ just as in [5]. This would not be competitive with Schönhage-Strassen method [10, Theorem 2.13]. We consider its ${\mathcal{Q}}d\log(d)\log(\log(d))$ complexity as the state of the art in polynomial multiplication and use this upper bound whenever needed in this article, despite the existence of an algorithm by Harvey and van der Hoeven [22] that multiplies polynomials with coefficients in a finite field with heuristic complexity ${\mathcal{Q}}d\log(d)$. Consequently, all the complexity estimates given in this article are fully unconditional.
Acknowledgements. We thank the anonymous referees for a careful reading of the manuscript and for helpful suggestions. We are also grateful to Grégoire Lecerf for valuable comments.
@@ 2. Summary and organization of the article
In this article, we consider specific evaluation and interpolation problems in genus one. We consider an elliptic curve $E$ over a field $\mathbf{K}$} of odd characteristic. We let $d$ be a power of two and we assume that there is a point $t$ of order $d$ in $E({\mathbf{K}})$. We let
@@
be the degree $d$ divisor associated to the group $\langle t\rangle$} generated by $t$. We take $b$ to be a $\mathbf{K}$}-rational point on $E$ such that $db$ is not equal to the point at infinity. We set
@@
A first application, presented in Section 13, is the construction of normal bases for finite field extensions of degree a power of two, allowing multiplication in time ${\mathcal{Q}}d\log(d)$. We build on the construction in [15] of an equivariant version of Chudnovsky’s interpolation method [12], and we use our complexity estimates for evaluation, interpolation and reduction in this context.
In Section 14, we construct Goppa [^2] In this article, we use the term Goppa code for any code obtained by evaluation of algebraic functions on a given algebraic curve at a collection of points on that curve. codes in genus one obtained by restricting the evaluation map $\mathop{\rm{ev}}\nolimits$ to the subspace of functions that are invariant by the elliptic involution $P\mapsto-P$ on $E$. The resulting code is a genus zero Goppa code. The advantage of seeing it as a genus one code is that we benefit from the acceleration due to the presence of the automorphism group $\langle t\rangle$}. This produces MDS $[d,d/2,d/2+1]$-codes that can be encoded and checked in time ${\mathcal{Q}}d\log(d)$ and decoded up to half the minimum distance at the expense of ${\mathcal{Q}}d\log^{2}(d)\log(\log(d))$ operations and comparisons in $\mathbf{K}$}. The only constraint for the existence of such codes is that the size of $\mathbf{K}$} be large enough with respect to $d$. This extends Lacan and Soro’s result [35] for Fermat fields.
@@ 3. Algorithms and data structures
For ease of reading, our pseudo-code descriptions of these algorithms are given in a recursive form. However, like the classical butterfly formulation of FFT algorithms à la Cooley–Tukey, they can be very easily unrolled to yield iterative functions. In fact, our algorithms are all straight-line programs.
@@ 4. Elliptic functions
We recall a few formulae from [15] regarding low-degree functions on elliptic curves. We let $\mathbf{K}$} be a field and $E$ the elliptic curve over $\mathbf{K}$} defined by the Weierstrass equation
@@ 5. Bases for elliptic function field extensions
Let $\mathbf{K}$} be a field with odd characteristic and let $E$ be an elliptic curve over $\mathbf{K}$}. Let $t$ be a point of order $d=2^{\delta}$ in $E({\mathbf{K}})$. We assume that $\delta\geqslant 1$. The group $\langle t\rangle$} generated by $t$ can be seen as an effective divisor of degree $d$ on $E$. We let ${\mathcal{L}}({\langle t\rangle})$ be the associated $\mathbf{K}$}-linear space. We construct two bases for ${\mathcal{L}}({\langle t\rangle})$ consisting of functions of degree $\leqslant 2$. We start with the following lemma from [15, Lemma 4].
@@
In what follows, we denote by UtoV_BaseChange and VtoU_BaseChange the subroutines that implement, respectively, the base change from $(u_{l})_{l\in{\mathbf{Z}}/d{\mathbf{Z}}}$ and $(v_{l})_{l\in{\mathbf{Z}}/d{\mathbf{Z}}}$ and vice versa. From Equations (6) and (7), we deduce that these routines require no more than ${\mathcal{Q}}d$ additions and multiplications by a constant in $\mathbf{K}$}.
6. Cyclic bidiagonal linear systems
We now consider linear sparse systems with a very special structure, called cyclic bidiagonal. These systems arise naturally in Section 8, and the ability to invert them efficiently is an essential ingredient for Section 9.
The determinant of $M$ is
$\det(M)=\prod_{i=0}^{d-1}b_{i}-(-1)^{d}\prod_{i=0}^{d-1}c_{i}.$
Thus, both evaluating and inverting a cyclic bidiagonal system of dimension $d$ require only $\mathcal{O}(d)$ operations. In the sequel, we denote by BiDiagonal_Evaluate and BiDiagonal_Invert the routines that implement, respectively, the direct matrix–vector product and this elimination procedure.
@@ 7. A Vélu isogeny of degree two
the point of order two in the group $\langle t\rangle$}. Following Vélu [38, 39] we write $T=(x(T),y(T))$ and set
@@ 8. Evaluation using elliptic butterflies
Under the hypotheses and notation of Sections 5 and 7, we let $b$ be a point in $E({\mathbf{K}})$ such that
$db\not=O.$
@@
$\hat{\varphi}(d^{\prime}b^{\prime})\not=O$
@@
There exists a constant $\mathcal{Q}$} such that the following is true. Let $\mathbf{K}$} be a field with odd characteristic. Let $E$ be an elliptic curve over $\mathbf{K}$}. Let $\delta\geqslant 1$ be an integer. Let $t$ be a point of order $d=2^{\delta}$ in $E({\mathbf{K}})$. Let $b$ be a point in $E({\mathbf{K}})$ such that $db\not=O$. There exists a straight-line program that on input a function $f$ in ${\mathcal{L}}({\langle t\rangle})$ given by its coordinates in either basis $u$ or $v$ defined in Section 5, computes the values $f(b+lt)$ for $l\in[0,d-1]$ at the expense of ${\mathcal{Q}}d\log(d)$ operations in $\mathbf{K}$}.
This results from the discussion above. For clarity, the evaluation procedure is presented as concise pseudo-code in Figure 2. $\square$
@@ 9. Interpolation
According to Equation (13) these $d^{\prime}$ scalars are the coordinates of $f^{1}=\theta\cdot f^{-}$ in the basis of ${\mathcal{L}}({\langle t^{\prime}\rangle}+O^{\prime}-U^{\prime})$ made of the $\theta\cdot(u_{l}-u_{l+d^{\prime}})$ for $l\in[0,d^{\prime}-1]$. We will use an auxiliary function $\xi_{b}$ on $E^{\prime}$ having degree $d^{\prime}+1$, polar divisor ${\langle t^{\prime}\rangle}+O^{\prime}$, and vanishing at $b^{\prime}+lt^{\prime}$ for every $l\in[0,d^{\prime}-1]$. Such a function is unique up to a multiplicative constant, and it has no pole and no zero at $U^{\prime}$ because $db\not=O$. We therefore assume that $\xi_{b}(U^{\prime})=1$. There exists $d^{\prime}+1$ scalars ${\mathfrak{d}}_{0}$, ${\mathfrak{d}}_{1}$, …, ${\mathfrak{d}}_{d^{\prime}-1}$, $\mathfrak{d}$$_{\ast}$ such that
@@
$f^{\star}=f^{\star}(U^{\prime})+\sum_{1\leqslant l\leqslant d^{\prime}-1}f^{\star}_{l}\cdot(v^{\prime}_{l}-v^{\prime}_{l}(U^{\prime}))\,.$
@@
Solving this system requires ${\mathcal{Q}}d$ operations in $\mathbf{K}$} as recalled in section 6. In the special case when $d^{\prime}=1$ we simply invoke Equation (17).
@@
There exists a constant $\mathcal{Q}$} such that the following is true. Let $\mathbf{K}$} be a field with odd characteristic. Let $E$ be an elliptic curve over $\mathbf{K}$}. Let $\delta\geqslant 1$ be an integer. Let $t$ be a point of order $d=2^{\delta}$ in $E({\mathbf{K}})$. Let $b$ be a point in $E({\mathbf{K}})$ such that $db\not=O$. There exists a straight-line program that on input $d$ scalars $(\alpha_{l})_{0\leqslant l\leqslant d-1}$, computes a function $f$ in ${\mathcal{L}}({\langle t\rangle})$, given by its coordinates in either basis $u$ or $v$ defined in Section 5, such that $f(b+lt)=\alpha_{l}$ for every $l\in[0,d-1]$, at the expense of ${\mathcal{Q}}d\log(d)$ operations in $\mathbf{K}$}.
This results from the discussion above. To aid understanding, the interpolation procedure is shown in compact pseudo-code in Figure 3. $\square$